# A Survey on Heterogeneous Computing Using SmartNICs and Emerging Data Processing Units (Expanded Preprint)

### v1

@@ 6.2 DPUs in Commercial and Infrastructure Applications
Multiple Tenants, Virtualization, and SDN: Applications running on SmartNICs are currently limited regarding tenants, because they require a level of mutual performance isolation which is unsupported by current software [30]; when the available resources include ASIC hardware accelerators in addition to the CPU, this complicates potential sharing mechanisms. ‘FairNIC’, created by Grant et al. [30], provides this isolation between tenants, enabling multiple applications to co-habitate with fair resource sharing on a single NIC without impacting each other’s performance, and demonstrating that sharing SmartNICs among virtual tenants is feasible, though it will require appropriate security mechanisms in the future. We also note in this section Shan Yizhou et al.’s [74] work on SuperNIC, also towards multi-tenancy and virtualization, which we have described previously. Partially related, Njavro et al. [66] apply DPU offloading to container overlay networks to decrease congestion in the host.
@@
In an older work, Yanfang et al. [52] put forward an architecture, named UNO, to provide Software-Defined Networking (SDN) controlled network function offloading through a virtual management plane, using multiple switches within the host, without interfering with data-center-wide data- and control-planes. The experimental findings reduce control-plane overhead, and demonstrate a potential reduction in power consumption by a factor of 2.
@@
Security: Mounir El Kiraf et al. [23] conducted a study in which they investigated the potential of DPUs as a scalable and transparent security solution. This solution incorporates Intrusion Detection Systems / Intrusion Prevention Systems (IDS/IPS) and supports multiple VPN connections. The study primarily focused on two aspects: first, the ability of the DPU to process, filter, and route all data to and from the operating system, thus reducing its workload and enhancing security; second, the scalability of the solution. The experimental results demonstrated the BlueField-1 DPU effectively ensures a transparent and scalable security solution in terms of VPN connections. Also, Sebastiano Miano et al. [62] directs the reader through the process of using a SmartNIC to offload mitigation of Distributed Denial of Service (DDoS) attacks, taking advantage of on-NIC hardware filtering.
@@
Kernel: Houxiang Ji et al. [40] proposed STYX, a framework that offloads memory optimization features from the Linux kernel to SmartNICs. It reduces the disruption of application execution, improving data center efficiency. STYX utilizes the SmartNIC’s RDMA capability to copy memory regions, and its compute capability to perform intensive operations, resulting in a significant decrease (55-89%) in 99th-percentile application latency[40].
HPC & Scientific Applications: DPUs have also been experimented with in the realm of high-performance computing (HPC) and scientific simulations - in particular, for tasks which require heavy communication patterns, such as halo exchanges, as found in molecular dynamics (MD) and particle simulations. In one such example, Sara Karamati et al. [43, 44] evaluated the benefits of using NVIDIA BlueField-2 DPUs as compute accelerators for MiniMD molecular dynamics simulations (proxy for LAMMPS). The BlueField-2 provided up to 20% speedup, with no real loss of simulation accuracy. However, perhaps of most interest with this work, they demonstrate the value of re-thinking current algorithms, highlighting the necessity of rearranging some algorithmic patterns to better take advantage of parallelism and DPUs’ strengths and communication patterns, and provide valuable advice thereon. Additionally, Matea Turalija et al. [86] also proposed DPUs for innovative methods of MD simulation acceleration, inspired by the Anton supercomputer because of its specially designed ASICS, and how DPUs also allow access to networking ASICs, which can be used to accelerate MD simulations in similar ways. This opens the door for other systems to potentially reach the same microsecond timescale needed for many real applications.
In a similar vein, Ulmer et al. [87] extend data services onto SmartNICs in order to leverage resource isolation, also for the benefit of HPC applications, such as particle simulations. The paper addresses how to construct software for implementing services on SmartNICs, and seeks to learn how useful they will be. Experimental results from a 100-node cluster using BlueField-2s indicate utility in data management tasks, though they are of course less performant than their hosts. Likewise, Liu et al. [58], described previously, perform experiments regarding particle data.
Finally, in the realm of geospatial computing, Kaymak et al. [46] experiment with offloading the ‘filter’ step in spatial join operations, or polygon intersection queries, leaving the heavier ‘refine’ step computations to the host devices. They use a cluster containing a number of BlueField DPUs for their experiment, and show that the offload provides a performance improvement.
@@
Blockchain: While the use of blockchain technology has become widespread across various industries, it is widely acknowledged that blockchains have the potential to have a significant environmental impact due to their high energy consumption and hardware requirements [42]. As a partial solution, Eish Kapoor et al. [42], building on their previous work [39], presented a blockchain infrastructure that runs completely on SmartNICs, called ‘BlockNIC’. Their findings point toward a somewhat more ecologically conscious way of building blockchain systems into future, under-utilized infrastructure [39].
DPU Management: Wang et al. [101] introduced a firmware upgrade system based on WAN to facilitate the secure over-the-air upgrading of DPUs, reducing operation and maintenance expenses. Empirical findings have indicated a nearly perfect success rate for DPU upgrades, alongside exceptional performance, in their particular application of simulating a Marine Engine Room.
@@ 6.3 DPUs in AI and ML Applications
Due to the current importance of artificial intelligence (AI) and machine learning (ML) in both research and industry, it is of particular interest that we note how DPUs have been used in conjunction with these technologies. Often, this means taking advantage of FPGAs to create hardware implementations of neural network architectures or operations, which can sometimes boast over two orders of magnitude faster inference [25, 27, 105], though it can also simply be done via computational offloading. See Table 6 for a breakdown of the types of neural networks etc. the research is exploring.
@@
        Deep Learning [105, 38, 37, 40]
        Convolutional NN (Non-SmartNIC: [71, 106, 22])
@@
In-Hardware / FPGA Impl. [105, 31, 25], (Non-SmartNIC: [71, 106, 22])
@@
Making Deep Learning (DL) Scalable: Anqi Guo et al. [31] proposed a framework called ‘FCsN’ for performing neural network inference on FPGA-based SmartNICs, aiming in general to improve the performance of HPC and data center processing. Their experimental results, tested on both DNNs and Graph Neural Networks (GNN), boast up to 10x speedups compared with a baseline of using MPI on CPUs. Later, Anqi Guo et al. [32] continue their work, proposing a heterogeneous SmartNIC system with both hardware and software co-designed, for distributed deep learning recommendation models (DLRMs). Their intent is to resolve the all-to-all communication bottleneck these models suffer from, and to increase the scalability of these models. Their approach enhances locality and computational efficiency, achieving a 2.1× inference latency speedup, and a 1.6× training throughput speedup. Both these works measure how their contributions affect the scalability of such systems.
@@
Omar Zawawi [100] identifies a “data stall” issue, where extensive data pre-processing at the CPU during DL training causes host devices to struggle to keep up with the training GPUs, causing the whole pipeline to stall and GPUs to spend time idle. He proposes offloading some of this pre-processing to the DPU in order to help mitigate the problem, as well as to improve energy efficiency within the data center by nature of the lighter-weight DPU energy requirements. On a similar note, Maroun Tork et al. [85] presented ‘Lynx’, a new network server architecture for offloading both data and control planes of network tasks to SmartNICs. The design is accelerator-centric, capable of running without CPU help by enabling more direct networking to the GPUs with lightweight I/O mechanisms. This allows for network servers to be hardware-accelerated, and attains greater than 4x throughput for their example GPU task of face-verification, though only 25% increase for an accelerated inference task, compared to host-centric architectures. However, [81] questions the efficacy of running Lynx on a SmartNIC versus the host.
@@
AI/ML for Network Traffic Management: Benjamin Fuhrer et al. [25] introduce a lightweight congestion control solution for datacenters using a reinforcement learning algorithm, RL-CC, which is converted into decision trees to achieve a 500x reduction in inference time. Deployment on NVIDIA ConnectX-6Dx SmartNICs demonstrated superior performance in managing bandwidth, latency, and packet loss in a balanced way, compared to existing algorithms like DCQCN and Swift across various benchmarks. On a related note, Hardware-based, on-NIC Binary Neural Networks (BNNs) are introduced to the data plane of FPGA SmartNICs by Siracusano et al. [27], in order to do classification and analysis of network traffic at a line rate of 40Gbps using machine learning. This hardware-accelerated neural network drastically improves latency compared to software implementations of similar networks (by up to two orders of magnitude), and appears to achieve decent accuracy, while freeing up the host’s CPU for other work. Their work is proof of the value of researching hardware implementations of neural networks.
@@
As for software implementations, Kasim Tasdemir et al. [83] investigate deploying Machine Learning (ML) algorithms on BlueField-3 DPUs for SQL injection detection. They test 20 different ML models on an SQL dataset, achieving near-real-time detection with a Passive Aggressive Classifier, with an accuracy of 99.78%.
@@
Co-Compute Offloading: Arpan Jain [37, 38] conducted a study to investigate the potential benefits of using the ARM cores of BlueField-2 DPUs to expedite the training of DNN models by offloading phases of the training to them; their study may represent the first such attempt. MPI is employed despite the heterogeneity of devices, and the researchers’ approach ultimately demonstrates up to 17.5% improvement in training duration on cutting-edge HPC clusters. They try several designs, and test them on multiple DL model types. They also measure scalability. Likewise, Shibahara et al. [78] point out the heavy network processing load for Federated Learning schemes, which require aggregating local weights from distributed learners and then sending global weights back out. They make use of the DPU’s cores to offload this aggregation, with a 1.39x speedup.
@@
In another offloading effort, Marina Perea-Trigo et al. [69] conducted a study to investigate the potential advantages of employing a DPU in order to mitigate the burden experienced by a continuously operating server. They proceeded to present the findings of their research, which involved quantifying the extent to which a CCTV system for weapon detection can alleviate workload, and evaluating its performance across various scenarios. As a result, they observed a remarkable reduction in workload of 43,123% over a 24-hour period, with savings exceeding 98% during nighttime intervals [69].
@@
Also, Tootaghaj et al. [84] proposed a new architecture, ‘Spike-Offload’, that strategically offloads workload spikes from microservices (important for edge ML workloads), when the SmartNIC has available compute power, reducing service level agreement (SLA) violations and vaunting better performance, lower energy consumption, and 40% potential reduction in capital expenditure.
Hardware Convolutional Neural Networks (CNN): Three works of note dealing with FPGA-based CNNs are those of Saidi et al., Jiang Zhu et al., and Zelin Du et al. [71, 106, 22]. However, it must be noted in our current context that these experiments, though effective, and relevant because their work may be somewhat transferrable to SmartNICs, were not performed on actual SmartNICs. Rather, these used ”Deep-Learning Processing Units” (which are not network cards), or a type of Xilinx FPGAs which are often included in SmartNICs.
Optical-Electronic Technologies: Zhong et al. [105] proposed ‘Lightning’, a pioneer in reconfigurable SmartNICs with photonic computing cores capable of doing DNN inference in real-time, a feat that outpaces NVIDIA A100 GPUs by over 300x, while using 300x less power, with similarly strong results for A100X DPUs and Brainwave SmartNICs. Their hardware-based approach avoids creating bottlenecks in packet processing and data movement. Their work demonstrates reasonably accurate photonic computation operations.
7 Overview Comments
@@
First, this is a very active field of research, with a large number of performance improvements being achievable with SmartNICs, in almost any type of work offloading. However, we note that many of these potential solutions are aimed at eating up the spare processing power and other resources of the DPU, meaning that they are not simultaneously applicable. Whether they are or are not, we recommend that the members of the research community be attentive to means of compatibly combining their methods with others’, in order to maximize the potential impact of their work and broaden its scope beyond their particular use-cases. Where methods are simultaneously applicable, such collaboration has the potential to result in impressively compounding improvements.
Second, it is clear from the research that the most impressive speedups and performance improvements shown are usually those which come from taking advantage of hardware acceleration; be it by properly making use of the on-board ASICs, by specialized FPGA programming, or by the advent of photonic computation, these continue to be the results which most often make orders of magnitude of difference, compared to current solutions. We recommend searching for every possible opportunity to apply the built-in accelerators, at least, to your target task. Other offloading is, of course, still worthwhile, and likely energy-saving, though it is likely capped by the computational capacity of the DPU, and the effects of reducing cache pollution and CPU congestion.
8 Conclusion
Much research is currently being done in exploring DPUs and SmartNICs as co-processors and accelerators, rather than simply as data center infrastructure offload and management devices, and this research is yielding favorable results. Therefore, interest in these heterogeneous computing systems has even spread to diverse fields outside of computer science and IT, such as the natural sciences and engineering, among others. New frameworks are being developed to provide the abstractions needed to bring these devices into wider usage, making them increasingly useful and effective. Despite this progress, several obstacles persist, but we believe that future research will yield resolutions to these challenges, and the niche for DPUs will continue to both develop and evolve.

### v2

@@ 6.1 DPUs in Commercial and Infrastructure Applications
Multiple Tenants, Virtualization, and SDN: Applications running on SmartNICs have been limited regarding tenants, because they require a level of mutual performance isolation which is unsupported by current software [26]; when the available resources include ASIC hardware accelerators in addition to the CPU, this complicates potential sharing mechanisms. ‘FairNIC’, created by Grant et al. [26], provides this isolation between tenants, enabling multiple applications to co-habitate with fair resource sharing on a single NIC without impacting each other’s performance, and demonstrating that sharing SmartNICs among virtual tenants is feasible, though it will require appropriate security mechanisms. ‘DORM’ attempts to improve upon such works, using optimization schemes to load-balance between the SmartNIC and host. We also note Yizhou et al.’s [83] work on ‘SuperNIC’, also towards multi-tenancy and virtualization, which we have described previously. Partially related, Njavro et al. [22] apply DPU offloading to container overlay networks to decrease congestion in the host.
@@
In an older work, Yanfang et al. [101] put forward an architecture, named UNO, to provide SDN-controlled network function offloading through a virtual management plane, using multiple switches within the host, without interfering with data-center-wide data– and control-planes. The experimental findings reduce control-plane overhead, and demonstrate a potential reduction in power consumption by a factor of 2.
@@
Security: Kiraf et al. [96] investigate the potential of DPUs as a scalable and transparent security solution. This solution incorporates Intrusion Detection Systems / Intrusion Prevention Systems (IDS/IPS) and supports multiple VPN connections. The study primarily focuses on two aspects: first, the ability of the DPU to process, filter, and route all data to and from the OS, thus reducing its workload and enhancing security; second, the scalability of the solution. The experimental results demonstrate effective use of the BlueField-1 DPU to ensure a transparent and scalable security solution in terms of VPN connections. Also, Miano et al. [49] direct the reader through the process of using a SmartNIC to offload mitigation of Distributed Denial of Service (DDoS) attacks, taking advantage of on-NIC hardware filtering.
@@
Kernel: Ji et al. [45] proposed STYX, a framework that offloads memory optimization features from the Linux kernel to SmartNICs. It reduces the disruption of application execution, improving data center efficiency. STYX utilizes the SmartNIC’s RDMA capability to copy memory regions, resulting in a 55-89% decrease in 99th-percentile application latency[45].
@@
Blockchain: While the use of blockchain technology has become widespread across various industries, it is widely acknowledged that blockchains have significant potential environmental impact due to their high energy consumption and hardware requirements [31]. In a slight reduction of this impact, Kapoor et al. [31], building on their previous work [102], presented a blockchain infrastructure that runs completely on SmartNICs to make use of spare processing power, called ‘BlockNIC’.
6.2 DPUs in HPC and Scientific Applications
In the realm of high-performance computing (HPC) and scientific simulations, a number of works have begun experimenting with using DPUs both to improve HPC systems, and to offload calculations, such as with large simulations—in particular, for tasks which require heavy communication patterns, such as ‘halo exchanges’, as found in molecular dynamics (MD) and particle simulations.
However, for such offloads to be effective, one often must start thinking outside the box; many algorithms are designed for certain architectures, and so it may often be beneficial to actually restructure current algorithms.
Molecular Dynamics Simulations: In one prime example of algorithmic restructuring, Karamati et al. [30] evaluated the benefits of using NVIDIA BlueField-2 DPUs as compute accelerators for MiniMD molecular dynamics simulations (a proxy for LAMMPS software). They redesign an existing algorithm to perform pipeline parallelism, which highlights the necessity of rearranging some algorithmic patterns to better take advantage of parallelism and DPUs’ strengths and communication patterns. They provide valuable advice to this end. With their restructuring, the BlueField-2 provided up to 20% speedup, with no real loss of simulation accuracy. Additionally, Turalija et al. [42] also proposed DPUs for innovative methods of MD simulation acceleration, inspired by the Anton supercomputer’s specially designed ASICs, since DPUs also allow access to networking ASICs, which can be used to accelerate MD simulations in similar ways to Anton. This opens the door for other systems to potentially reach the same microsecond timescale needed for many real-world applications.
In a similar vein, Ulmer et al. [14] extend data services onto SmartNICs in order to leverage resource isolation, also for the benefit of HPC applications, such as particle simulations. The paper addresses how to construct software for implementing services on SmartNICs, and seeks to learn how useful they will be. Experimental results from a 100-node cluster using BlueField-2s indicate utility in data management tasks, though they are, as expected, less performant than their hosts. Likewise, Liu et al. [18], described previously, perform experiments regarding particle data. Both these works use Apache Arrow.
HPC Systems: Horany et al.’s work [59] aims to help with allocating network-wide resources for large-scale systems. Tong et al.’s work [93] on CPU-DPU hybrid learned indexes also targets the needs of HPC systems, and Suresh et al. [56] lay the groundwork for offloading vector and matrix operations and test it on Krylov solvers.
Geospatial Computing: Kaymak et al. [58] experiment with offloading the ‘filter’ step in spatial join operations (polygon intersection queries), leaving the heavier ‘refine’ step computations to the host devices. They use a cluster of DPUs for their experiment, showing some performance improvement. They also briefly measure scalability.
@@ 6.3 DPUs in AI and ML Applications
Due to the current importance of artificial intelligence (AI) and machine learning (ML) in both research and industry, it is of particular interest that we note how DPUs have been used in conjunction with these technologies. Often, this means taking advantage of FPGAs to create hardware implementations of neural network architectures or operations, which can sometimes boast over two orders of magnitude faster inference [82, 98, 77], although it can also simply be done via computational offloading. See Table 7 for a breakdown of the types of neural networks etc. that the research is exploring.
@@
     Deep Learning [77, 55, 45]
@@
In-Hardware / FPGA Impl. [77, 97, 82]
@@
Making Deep Learning (DL) Scalable: There are various problems with scaling DL systems, which may be addressed via clever use of SmartNICs. Guo et al. [97] proposed a framework called ‘FCsN’ for performing neural network inference on FPGA-based SmartNICs, aiming to improve the performance of HPC and data center processing. Their experimental results, tested on both DNNs and Graph Neural Networks (GNN), boast up to 10x speedups compared with a baseline of using MPI on CPUs. Later, Guo et al. [76] continue their work, proposing a heterogeneous SmartNIC system with both hardware and software co-designed, for distributed deep learning recommendation models (DLRMs). Their intent is to resolve the all-to-all communication bottleneck these models suffer from and increase their scalability. Their approach enhances locality and computational efficiency, achieving a 2.1× inference latency speedup, and a 1.6× training throughput speedup. Both these works measure how their contributions affect scalability.
@@
Zawawi [32] identifies a “data stall” issue, where extensive data preprocessing at the CPU during DL training causes host devices to struggle to keep up with the training GPUs, causing the whole pipeline to stall and GPUs to spend time idle. They help mitigate this by offloading some of this preprocessing to the DPU, also somewhat improving energy efficiency within the data center by nature of the lighter-weight DPU energy requirements. On a similar note, Tork et al. [75] presented ‘Lynx’, a new network server architecture for offloading both data and control planes of network tasks to SmartNICs. The design is accelerator-centric, capable of running without CPU help by enabling more direct networking to the GPUs. This attains greater than 4x throughput for their example GPU task of face-verification, and 25% increase for an accelerated inference task, compared to host-centric architectures. However, [23] questions the efficacy of running Lynx on a SmartNIC versus the host.
@@
AI/ML for Network Traffic Management: Fuhrer et al. [82] introduce a lightweight congestion control solution for datacenters using a reinforcement learning algorithm, RL-CC, which is converted into decision trees to achieve a 500x reduction in inference time. Deployment on NVIDIA ConnectX-6 Dx SmartNICs demonstrated superior performance in managing bandwidth, latency, and packet loss in a balanced way, compared to existing algorithms like DCQCN and Swift, across various benchmarks. On a related note, Hardware-based, on-NIC Binary Neural Networks (BNNs) are introduced to the data plane of FPGA SmartNICs by Siracusano et al. [98], in order to do classification and analysis of network traffic at a line rate of 40Gbps using machine learning. This hardware-accelerated neural network drastically improves latency compared to software implementations of similar networks (by up to two orders of magnitude), and appears to achieve decent accuracy, while freeing up the host’s CPU for other work. Their work demonstrates significant value in researching hardware implementations of neural networks.
@@
As for software implementations, Tasdemir et al. [92] investigate deploying Machine Learning (ML) algorithms on BlueField-3 DPUs for SQL injection detection. They test 20 different ML models on an SQL dataset, achieving near-real-time detection with a Passive Aggressive Classifier, with an accuracy of 99.78%.
@@
Co-Compute Offloading: Arpan Jain [55] investigate the potential benefits of using the Arm cores of BlueField-2 DPUs to expedite the training of DNN models by offloading phases of the training to them; their study may represent the first such attempt. MPI is employed despite the heterogeneity of devices, and the researchers’ approach ultimately demonstrates up to 17.5% improvement in training duration on cutting-edge HPC clusters. They try several designs, and test them on multiple DL model types. They also measure scalability. Likewise, Shibahara et al. [50] point out the heavy network processing load for Federated Learning schemes, which require aggregating local weights from distributed learners and then sending global weights back out. They make use of the DPU’s cores to offload this aggregation, with a 1.39x speedup.
@@
In another offloading effort, Perea-Trigo et al. [94] investigate the potential advantages of employing an NVIDIA BlueField-2 DPU to mitigate the burden experienced by a continuously operating server. They quantify the extent to which a DPU can alleviate the workload of a CCTV system for weapon detection, evaluating its performance across various scenarios. They observed a remarkable reduction in workload of 43,123% over a 24-hour period, with savings exceeding 98% during nighttime intervals [94].
@@
Also, Tootaghaj et al. [95] proposed a new architecture, ‘Spike-Offload’, which strategically offloads workload spikes from microservices (important for edge ML workloads), when the SmartNIC has available compute power, thus reducing service level agreement (SLA) violations and giving better performance, lower energy consumption, and 40% potential reduction in capital expenditure.
Optical-Electronic Technologies: Z. Zhong et al. [77] proposed ‘Lightning’, a hardware-based prototype SmartNIC with photonic computing cores. Lightning is a pioneer in leveraging the fast data paths of a SmartNIC to overcome the bottleneck of controlling and moving data from the electronic domain to the photonic domain, where DNN inference can be performed much faster. To achieve this, they modify an algorithm to be more pipelined. Their prototype and simulations are capable of doing DNN inference in real-time, outpacing the same inference done purely on NVIDIA A100 GPUs by over 300x (and thus also using 300x less power).
7 Conclusion
@@
First, this is a very active field of research, and a large number of works demonstrate achievable performance improvements using SmartNICs, offloading a broad variety of types of work. In areas where there is potential for future compatibility, we recommend that the members of the research community pursue integration of their solutions, being thoughtful of eventual means of combining their methods with others’, in order to maximize the potential impact of their work and broaden its scope beyond their particular use-cases. Such collaboration likely has the potential to result in impressively compounding improvements. Even in cases where the same resource is maximized by two or more solutions, such as the many works which achieve performance boosts by filling the DPU processors to capacity, concepts from individual solutions might be combined into a single integrated solution. At the very least, the community would likely benefit from some form of collaborative, well-supported, diverse, cohesive suite of such solutions, each with its varied strengths, to accessibly choose from when developing with DPUs. Even among the SDKs provided by DPU vendors, standardization and cross-compatibility are currently lacking [1].
Second, it is clear from a large segment of the research that the largest speedups and performance improvements shown are almost exclusively those which come from taking advantage of hardware acceleration: making use of the on-board ASICs, specialized FPGA programming, laying the groundwork for photonic computation, etc. These hardware-leaning efforts continue to be the results which most often make orders of magnitude of difference, compared to current methods. We thus recommend searching for every possible opportunity to apply the built-in accelerators to your target task. Standard parallel computational offloading may still be worthwhile, and likely energy-saving, though it is usually capped by the computational capacity of the DPU and the effects of reducing cache pollution and CPU congestion in the host.
Finally, much research is currently being done in exploring DPUs and SmartNICs as heterogeneous co-processors and accelerators, as opposed to simply as data center infrastructure offload and management devices, for nearly any kind of work-offloading or benchmarking. Much of this research appears to be yielding favorable performance improvements, ranging from meager to large. Interest in these heterogeneous computing systems has even spread to diverse fields outside of computer science and IT, such as the natural sciences and engineering. Because such on-NIC computational applications are not the primary target of most SmartNICs, these research angles provide momentum to the field in an interesting new direction; likewise, we see the development of DPUs with higher and higher compute capacity mirroring (or perhaps causing) this trend. New frameworks are being developed to provide the abstractions needed to bring these devices into wider usage, making them increasingly accessible and effective. Despite this progress, several obstacles persist, but researchers appear to be working to address these challenges as the niche for DPUs continues to develop and evolve.